OpenStudy (anonymous):

please help! Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (anonymous):

The correct answer is "B". From the given equation [y=3X + 2] we know that its slope is 3...parallel lines have the same slope. So now you have this much of the new equation you are to find...[y=3x + ...] . Now you need to find "b" from the Y Intercept Form, Y=mx+b. You can finf "b" by using the given data point (5,2) in your unfinished "new" equqtion--Y=3x + b and solve for "b" after insertng 5 for x and 2 for y. You should now have Y=3x+b 2=3(5)+b multiply 3x5, then subtract (-) 15 from the right side of the =, and then subtrct it also from thr left side of the = ... -15 +2=15 -15 +b -13 = b So b= -13, now go back to you unfinished "new" equation and plug in "b"... The equqtion you were to find is Y=3x-13.